# Parallel Virutal Machine PICL Portable Instrumented Communication Library

### 13 Mar 2026 v1

@@ Abstract
This paper presents NCCL EP (Expert Parallelism), a ground-up MoE communication library built entirely on NCCL’s Device API. NCCL EP provides unified ncclEpDispatch and ncclEpCombine primitives with both C and Python interfaces, supporting Low-Latency (LL) mode for inference decoding and High-Throughput (HT) mode for training and inference prefill. LL targets small batch sizes (1–128 tokens) using direct all-to-all RDMA+NVLink mesh connectivity with double-buffered communication for overlapping dispatch and combine phases. HT targets large batches (4096+ tokens) using hierarchical communication that aggregates tokens within NVLink domains before inter-node RDMA transmission. Both modes leverage Device API for both intra- and inter-node communications, taking advantage of its topology awareness and optimized GPU-initiated implementation.
We evaluate NCCL EP on an H100-based cluster across multi-node configurations, demonstrating competitive LL kernel performance and presenting end-to-end results with vLLM integration. By building MoE communication natively within NCCL, NCCL EP provides a supported path for expert parallelism on current and emerging NVIDIA platforms.
@@ I Introduction
Low-latency (LL) Mode for inference decoding. The design is derived from DeepEP [deepep2025] low-latency kernel with memory footprint optimizations. The implementation employs full all-to-all mesh connectivity with double-buffered communication.
@@ I-A Contributions
    Device-Initiated Communication. Both LL and HT rely on GPU-initiated communication implemented via NCCL Device API. Excluding the CPU from the data path enables asynchronous RDMA communication.
@@
    Optimized LL design. We propose a memory-optimized version of the LL kernels that achieves an order-of-magnitude reduction in memory footprint.
To evaluate NCCL EP in a production inference serving system, we provide its integration with vLLM and Megatron-LM frameworks. Building on top of that, we conduct comprehensive evaluation against DeepEP across 1–8 node configurations on an H100-based cluster, demonstrating competitive LL kernel performance and presenting end-to-end results with vLLM integration.
@@ II-D The MoE Communication Landscape
Library Unified API LL Mode HT Mode C API Python Inter-node Target Workload
@@ III-B Core Operations
ncclEpComplete(handle, config, stream) Finalize staged operation (LL mode). Blocks until receives complete.
@@ III-C1 MoE Group
The ncclEpGroup_t represents the top-level resource container, created via ncclEpCreateGroup() from an existing NCCL communicator. The group configuration is determined by the algorithm mode (LL or HT), number of experts, maximum tokens per rank, and buffer-sizing parameters. The group encapsulates:
@@ III-C2 MoE Handle
The ncclEpHandle_t maintains per-pass state, created via ncclEpCreateHandle() with the routing information (topk_idx) that specifies which experts each token should be dispatched to. In HT mode, handle creation triggers metadata exchange across ranks to compute token distribution; in LL mode, this exchange occurs implicitly during dispatch.
@@ III-D Algorithm Modes
LL Mode is optimized for low latency. It takes token data as input and produces a three-dimensional, expert-major output tensor ready to be consumed by subsequent expert-defined GEMM operations (Figure 3). LL execution can be staged (send_only=1) enabling expert computation to overlap with data transfer. Additionally, LL supports overlapping consecutive dispatch and combine phases through double buffering.
@@
TABLE III: Comparison of HT and LL kernel modes
@@ III-E3 Tensor Layouts
  • LL mode (Figure 3): 3D with shape [num_local_experts $\times$ tokens_per_expert $\times$ hidden], where tokens are groupped by receiving expert. This expert-major layout enables direct input to grouped GEMM operations.
@@
Combine: Input shapes match the corresponding dispatch outputs (3D for LL mode as in Figure 3, 2D for HT mode as in Figure 4); output shapes match dispatch inputs, restoring tokens to their original order.
III-F Usage Patterns: LL and HT in Training and Inference
@@ III-F1 Training with HT Mode
III-F2 Inference with LL Mode
For decoding, LL mode with staged execution enables overlapping multiple micro-batches:
@@ IV Low-Latency Kernels
NCCL EP LL mode targets the decode stage of inference, where small batches (typically 1–128 tokens) make end-to-end latency the primary concern. DeepEP [deepep2025] is the state of the art implementation, that leverages device-initiated communication, emphasizing dispatch-combine time reduction and computation-communication overlap. NCCL EP adopts DeepEP design, replacing the the communication backend with NCCL Device API, and applying memory footprint and communication optimizations.
@@ IV-A MoE Model
NCCL EP assumes an block-wise distribution of $E$ experts by $N$ application ranks, with $L=\lceil E/N\rceil$ experts per rank. On each rank $r$, the attention sub-layer produces a batch of $B$ fixed size tokens. For a token index $t\in[0,B)$ on rank $r$, the MoE router defines a vector $R(r,t)$ of $K$ expert indices to which $t$ is routed.
The logical communication occurs between two sides: data-parallel ranks (superscript $DP$) and experts ($E$). Every data-parallel rank $r$ communicates with all experts forming $E$ valid expert-rank pairs $\Gamma^{DP}_{r}=\{(e,r)|e\in[0,E)\}$, where $rem^{DP}(e,r)=\lfloor e/N\rfloor$ is the remote rank. Each expert $e$ communicates with all ranks. Thus, rank $r$ that hosts $L$ experts also forms $N\cdot L=E$ valid expert-rank pairs $\Gamma^{E}_{r}=\{(e,r^{\prime})|r^{\prime}\in[0,N),\lfloor e/N\rfloor=r\}$, with remote peer defined as $rem^{E}(e,r)=r$. The number of tokens communicated between pair $(e,r)$ is denoted as $m^{s}(e,r)$, $s~\in~\{R,E\}$.
@@ IV-B DeepEP Communication Model
LL kernels employ full NN-to-NN mesh connectivity. Every GPU (represented by a rank) directly communicates with any other GPU through the most efficient available channel, enabling the high routing flexibility required for dynamic expert distribution in MoE models.
@@
LL mode allocates two internal buffers used in double-buffering scheme to overlap dispatch and combine operations. Each buffer is organized into three regions: (a) coordination–per-$(e,r)$ pair signals, (b) send–a bounce buffer to assemble message payload, and (c) receive–for messages from remote ranks.
@@
Each sub-region is divided into $B$ payload slots. The slot size $P$ is determined by the operation’s per-token payload message. The LL dispatch payload includes (a) a header containing the source token index, (b) token data, and (c) optional FP8 quantization scales when precision reduction is enabled. The combine payload includes only token data. For DeepSeek v3 model, the token data size is 7168 B for FP8 data type and 14336 B for BF16; FP8 quantization scales contain 56 floats (224 B).
@@
After all $m^{s}(e,r)$ tokens corresponding to a pair $(e,r)$ were enqueued, an update-and-flush operation on $idx^{s}(e,r)$’th counter of rank $r^{\prime}$ with operation-specific value $V$ is issued. For dispatch, $V=(m^{s}(r,e)+1)$, communicating the number of valid slots in the respective sub-region (the additive of $1$ allows non-zero updates when $m^{s}(e,r)=0$). For combine, $V=1$ to flush tokens and signal $DP$ side.
@@ IV-C DeepEP Parallelization model
The LL kernels are executed on a set of $S$ SMs, each SM $i$ is sub-divided on $W=32$ warps $\{\omega_{ij},i\in[0,S),j\in[0,W)\}$.
@@ IV-D NCCL EP implementation
NCCL EP implements communication channels through NCCL Device API (Section II-C). When the destination is in the same NVLink domain (same LSA team), LL kernels leverage NVLink network using load/store instructions. Otherwise, the RDMA network is used via NCCL GIN API.
@@
The new scheme requires minimal changes in the LL implementation. As the coordination region footprint is low, the coordination logic was not changed. The footprint reduction is described by (3), and for $N=64$ ranks, $E=512$ experts, and top-k $K=8$ yields $\approx 14$ times improvement.
@@ V-D Comparison with Low-Latency Mode
Table III summarizes the key differences between HT and LL modes. The HT output is a 2D tensor with shape [num_recv_tokens $\times$ hidden], along with per-expert token counts indicating boundaries. This format suits grouped GEMM operations that process variable-sized expert batches. In contrast, LL mode produces a 3D expert-major layout optimized for the small batch sizes typical of decode.
@@ VI-B Megatron-LM Integration
vLLM implements expert-parallel MoE inference with pluggable All2All backends (naive, pplx, DeepEP’s HT/LL). NCCL EP integrates as an additional backend (nccl_high_throughput or nccl_low_latency). Unlike Megatron, vLLM supports both algorithm modes; tensor layouts follow the mode-specific formats described in Section III-E.
@@
Buffer and activation formats. The buffer provides per-expert token counts (device or host) to guide GEMM and combine, avoiding wasted work on padding. In LL mode, the batched-per-expert layout enables buffer preallocation for lower latency. In HT mode, the 2D concatenated layout with per-expert counts supports deterministic ordering.
@@ VII Performance Evaluation
This section evaluates NCCL EP performance against DeepEP, the widely-adopted MoE communication library. We present Low-Latency (LL) kernel results for inference-decode (Section VII-A) and vLLM integration benchmarks (Section VII-B). High-Throughput (HT) kernel evaluation and optimization is ongoing; HT results and Megatron-LM integration benchmarks will be released after these optimizations are complete.
@@ VII-A Low-Latency Kernels
LL kernels optimize for MoE inference-decode with small token batches, using full all-to-all RDMA mesh connectivity with per-expert signals and hybrid NVLink-RDMA paths. NCCL EP currently only supports the hybrid NVLink-RDMA path. We evaluate with 256 experts, hidden dimension 7168, 128 tokens, top-k 8 routing, and BF16 precision for both dispatch and combine operations.
@@
Figure 7: LL dispatch throughput: NCCL EP vs DeepEP.
Figure 8: LL combine throughput: NCCL EP vs DeepEP.
@@
Table VII compares end-to-end serving metrics for NCCL EP and DeepEP integrated with vLLM on the EOS cluster. Unlike the LL kernel evaluation in Section VII-A, which compared both implementations over the same NCCL transport, this evaluation runs DeepEP v1.2.1 with its default NVSHMEM 3.2.5/IBGDA backend—reflecting the configuration users deploy in production. Both backends use vLLM 0.10. We use 1-node (8 GPU), 2-node (16 GPU), and 4-node (32 GPU) configurations with the Qwen3-30B-A3B model [qwen3-technical-report], a 30B-parameter MoE LLM, serving 1000 requests with a maximum concurrency of 32. All configurations use TP=1 with DP=8, 16, and 32 (equal to EP) respectively. Each configuration was run 4 times per backend; reported values are means after removing statistical outliers via IQR filtering.
@@
NCCL EP trails DeepEP by 7–10% in throughput and 7–9% in ITL/TPOT across all scales. Prior work on NCCL GIN [ncclgin] showed that DeepEP over NCCL GIN and DeepEP over NVSHMEM achieve broadly comparable transport-level performance, with each favoring different operations and scales. The consistent ITL/TPOT gap is therefore not explained by the transport difference; the LL combine overhead observed in Section VII-A is a contributing factor. Investigating and optimizing this gap is ongoing for future NCCL EP releases.
@@ VIII-A Device-Initiated MoE Libraries
DeepEP [deepep2025] pioneered device-initiated MoE communication with NVSHMEM and IBGDA. It provides FP8 quantization, hook-based overlap, and optimizations for DeepSeek-V3’s group-limited gating. DeepEP exposes separate Python APIs for LL and HT modes. NCCL EP’s LL kernels adapt DeepEP’s design to NCCL GIN.
@@ VIII-D NCCL EP’s Positioning
    Unified API: DeepEP, HybridEP, and pplx-kernels require separate interfaces for LL and HT modes; NCCL EP offers unified ncclEpDispatch/ncclEpCombine primitives with the algorithm mode selected at group creation time.
@@
By combining proven kernel techniques from DeepEP (LL) and HybridEP (HT) with NCCL’s production infrastructure, NCCL EP delivers high-performance MoE communication within the standard GPU communication ecosystem.

### 24 Mar 2026 v2

@@ Abstract
This paper presents NCCL EP (Expert Parallelism), a ground-up MoE communication library built entirely on NCCL’s Device API. NCCL EP provides unified ncclEpDispatch and ncclEpCombine primitives with both C and Python interfaces, supporting Low-Latency ($L$) mode for inference decoding and High-Throughput (HT) mode for training and inference prefill. $L$ targets small batch sizes (1–128 tokens) using direct all-to-all RDMA+NVLink mesh connectivity with double-buffered communication for overlapping dispatch and combine phases. HT targets large batches (4096+ tokens) using hierarchical communication that aggregates tokens within NVLink domains before inter-node RDMA transmission. Both modes leverage Device API for both intra- and inter-node communications, taking advantage of its topology awareness and optimized GPU-initiated implementation.
We evaluate NCCL EP on an H100-based cluster across multi-node configurations, demonstrating competitive $L$ kernel performance and presenting end-to-end results with vLLM integration. By building MoE communication natively within NCCL, NCCL EP provides a supported path for expert parallelism on current and emerging NVIDIA platforms.
@@ I Introduction
Low-latency ($L$) Mode for inference decoding. The design is derived from DeepEP [deepep2025] low-latency kernel with memory footprint optimizations. The implementation employs full all-to-all mesh connectivity with double-buffered communication.
@@ I-A Contributions
    Device-Initiated Communication. Both $L$ and HT rely on GPU-initiated communication implemented via NCCL Device API. Excluding the CPU from the data path enables asynchronous RDMA communication.
@@
    Optimized $L$ design. We propose a memory-optimized version of the $L$ kernels that achieves an order-of-magnitude reduction in memory footprint.
To evaluate NCCL EP in a production inference serving system, we provide its integration with vLLM and Megatron-LM frameworks. Building on top of that, we conduct comprehensive evaluation against DeepEP across 1–8 node configurations on an H100-based cluster, demonstrating competitive $L$ kernel performance and presenting end-to-end results with vLLM integration.
@@ II-D The MoE Communication Landscape
Library Unified API $L$ Mode HT Mode C API Python Inter-node Target Workload
@@ III-B Core Operations
ncclEpComplete(handle, config, stream) Finalize staged operation ($L$ mode). Blocks until receives complete.
@@ III-C1 MoE Group
The ncclEpGroup_t represents the top-level resource container, created via ncclEpCreateGroup() from an existing NCCL communicator. The group configuration is determined by the algorithm mode ($L$ or HT), number of experts, maximum tokens per rank, and buffer-sizing parameters. The group encapsulates:
@@ III-C2 MoE Handle
The ncclEpHandle_t maintains per-pass state, created via ncclEpCreateHandle() with the routing information (topk_idx) that specifies which experts each token should be dispatched to. In HT mode, handle creation triggers metadata exchange across ranks to compute token distribution; in $L$ mode, this exchange occurs implicitly during dispatch.
@@ III-D Algorithm Modes
$L$ Mode is optimized for low latency. It takes token data as input and produces a three-dimensional, expert-major output tensor ready to be consumed by subsequent expert-defined GEMM operations (Figure 3). $L$ execution can be staged (send_only=1) enabling expert computation to overlap with data transfer. Additionally, $L$ supports overlapping consecutive dispatch and combine phases through double buffering.
@@
TABLE III: Comparison of HT and $L$ kernel modes
@@ III-E3 Tensor Layouts
  • $L$ mode (Figure 3): 3D with shape [num_local_experts $\times$ tokens_per_expert $\times$ hidden], where tokens are groupped by receiving expert. This expert-major layout enables direct input to grouped GEMM operations.
@@
Combine: Input shapes match the corresponding dispatch outputs (3D for $L$ mode as in Figure 3, 2D for HT mode as in Figure 4); output shapes match dispatch inputs, restoring tokens to their original order.
III-F Usage Patterns: $L$ and HT in Training and Inference
@@ III-F1 Training with HT Mode
III-F2 Inference with $L$ Mode
For decoding, $L$ mode with staged execution enables overlapping multiple micro-batches:
@@ IV Low-Latency Kernels
NCCL EP $L$ mode targets the decode stage of inference, where small batches (typically 1–128 tokens) make end-to-end latency the primary concern. DeepEP [deepep2025] is the state of the art implementation, that leverages device-initiated communication, emphasizing dispatch-combine time reduction and computation-communication overlap. NCCL EP adopts DeepEP design, replacing the the communication backend with NCCL Device API, and applying memory footprint and communication optimizations.
@@ IV-A MoE Model
NCCL EP assumes an block-wise distribution of $E$ experts by NN application ranks, with $L=\lceil E/N\rceil$ experts per rank. On each rank $r$, the attention sub-layer produces a batch of $B$ fixed size tokens. For a token index $t\in[0,B)$ on rank $r$, the MoE router defines a vector $R(r,t)$ of $K$ expert indices to which $t$ is routed.
The logical communication occurs between two sides: data-parallel ranks (superscript DPDP) and experts ($E$). Every data-parallel rank $r$ communicates with all experts forming $E$ valid expert-rank pairs $\Gamma^{DP}_{r}=\{(e,r)|e\in[0,E)\}$, where $rem^{DP}(e,r)=\lfloor e/N\rfloor$ is the remote rank. Each expert $e$ communicates with all ranks. Thus, rank $r$ that hosts $L$ experts also forms $N\cdot L=E$ valid expert-rank pairs $\Gamma^{E}_{r}=\{(e,r^{\prime})|r^{\prime}\in[0,N),\lfloor e/N\rfloor=r\}$, with remote peer defined as $rem^{E}(e,r)=r$. The number of tokens communicated between pair $(e,r)$ is denoted as $m^{s}(e,r)$, $s~\in~\{R,E\}$.
@@ IV-B DeepEP Communication Model
$L$ kernels employ full NN-to-NN mesh connectivity. Every GPU (represented by a rank) directly communicates with any other GPU through the most efficient available channel, enabling the high routing flexibility required for dynamic expert distribution in MoE models.
@@
$L$ mode allocates two internal buffers used in double-buffering scheme to overlap dispatch and combine operations. Each buffer is organized into three regions: (a) coordination–per-$(e,r)$ pair signals, (b) send–a bounce buffer to assemble message payload, and (c) receive–for messages from remote ranks.
@@
Each sub-region is divided into $B$ payload slots. The slot size $P$ is determined by the operation’s per-token payload message. The $L$ dispatch payload includes (a) a header containing the source token index, (b) token data, and (c) optional FP8 quantization scales when precision reduction is enabled. The combine payload includes only token data. For DeepSeek v3 model, the token data size is 7168 B for FP8 data type and 14336 B for BF16; FP8 quantization scales contain 56 floats (224 B).
@@
After all $m^{s}(e,r)$ tokens corresponding to a pair $(e,r)$ were enqueued, an update-and-flush operation on $idx^{s}(e,r)$’th counter of rank $r^{\prime}$ with operation-specific value $V$ is issued. For dispatch, $V=(m^{s}(r,e)+1)$, communicating the number of valid slots in the respective sub-region (the additive of 11 allows non-zero updates when $m^{s}(e,r)=0$). For combine, $V=1$ to flush tokens and signal DPDP side.
@@ IV-C DeepEP Parallelization model
The $L$ kernels are executed on a set of $S$ SMs, each SM $i$ is sub-divided on $W=32$ warps $\{\omega_{ij},i\in[0,S),j\in[0,W)\}$.
@@ IV-D NCCL EP implementation
NCCL EP implements communication channels through NCCL Device API (Section II-C). When the destination is in the same NVLink domain (same LSA team), $L$ kernels leverage NVLink network using load/store instructions. Otherwise, the RDMA network is used via NCCL GIN API.
@@
The new scheme requires minimal changes in the $L$ implementation. As the coordination region footprint is low, the coordination logic was not changed. The footprint reduction is described by (3), and for $N=64$ ranks, $E=512$ experts, and top-k $K=8$ yields $\approx 14$ times improvement.
@@ V-D Comparison with Low-Latency Mode
Table III summarizes the key differences between HT and $L$ modes. The HT output is a 2D tensor with shape [num_recv_tokens $\times$ hidden], along with per-expert token counts indicating boundaries. This format suits grouped GEMM operations that process variable-sized expert batches. In contrast, $L$ mode produces a 3D expert-major layout optimized for the small batch sizes typical of decode.
@@ VI-B Megatron-LM Integration
vLLM implements expert-parallel MoE inference with pluggable All2All backends (naive, pplx, DeepEP’s HT/$L$). NCCL EP integrates as an additional backend (nccl_high_throughput or nccl_low_latency). Unlike Megatron, vLLM supports both algorithm modes; tensor layouts follow the mode-specific formats described in Section III-E.
@@
Buffer and activation formats. The buffer provides per-expert token counts (device or host) to guide GEMM and combine, avoiding wasted work on padding. In $L$ mode, the batched-per-expert layout enables buffer preallocation for lower latency. In HT mode, the 2D concatenated layout with per-expert counts supports deterministic ordering.
@@ VII Performance Evaluation
This section evaluates NCCL EP performance against DeepEP, the widely-adopted MoE communication library. We present Low-Latency ($L$) kernel results for inference-decode (Section VII-A) and vLLM integration benchmarks (Section VII-B). High-Throughput (HT) kernel evaluation and optimization is ongoing; HT results and Megatron-LM integration benchmarks will be released after these optimizations are complete.
@@ VII-A Low-Latency Kernels
$L$ kernels optimize for MoE inference-decode with small token batches, using full all-to-all RDMA mesh connectivity with per-expert signals and hybrid NVLink-RDMA paths. NCCL EP currently only supports the hybrid NVLink-RDMA path. We evaluate with 256 experts, hidden dimension 7168, 128 tokens, top-k 8 routing, and BF16 precision for both dispatch and combine operations.
@@
Figure 7: $L$ dispatch throughput: NCCL EP vs DeepEP.
Figure 8: $L$ combine throughput: NCCL EP vs DeepEP.
@@
Table VII compares end-to-end serving metrics for NCCL EP and DeepEP integrated with vLLM on the EOS cluster. Unlike the $L$ kernel evaluation in Section VII-A, which compared both implementations over the same NCCL transport, this evaluation runs DeepEP v1.2.1 with its default NVSHMEM 3.2.5/IBGDA backend—reflecting the configuration users deploy in production. Both backends use vLLM 0.10. We use 1-node (8 GPU), 2-node (16 GPU), and 4-node (32 GPU) configurations with the Qwen3-30B-A3B model [qwen3-technical-report], a 30B-parameter MoE LLM, serving 1000 requests with a maximum concurrency of 32. All configurations use TP=1 with DP=8, 16, and 32 (equal to EP) respectively. Each configuration was run 4 times per backend; reported values are means after removing statistical outliers via IQR filtering.
@@
NCCL EP trails DeepEP by 7–10% in throughput and 7–9% in ITL/TPOT across all scales. Prior work on NCCL GIN [ncclgin] showed that DeepEP over NCCL GIN and DeepEP over NVSHMEM achieve broadly comparable transport-level performance, with each favoring different operations and scales. The consistent ITL/TPOT gap is therefore not explained by the transport difference; the $L$ combine overhead observed in Section VII-A is a contributing factor. Investigating and optimizing this gap is ongoing for future NCCL EP releases.
@@ VIII-A Device-Initiated MoE Libraries
DeepEP [deepep2025] pioneered device-initiated MoE communication with NVSHMEM and IBGDA. It provides FP8 quantization, hook-based overlap, and optimizations for DeepSeek-V3’s group-limited gating. DeepEP exposes separate Python APIs for $L$ and HT modes. NCCL EP’s $L$ kernels adapt DeepEP’s design to NCCL GIN.
@@ VIII-D NCCL EP’s Positioning
    Unified API: DeepEP, HybridEP, and pplx-kernels require separate interfaces for $L$ and HT modes; NCCL EP offers unified ncclEpDispatch/ncclEpCombine primitives with the algorithm mode selected at group creation time.
@@
By combining proven kernel techniques from DeepEP ($L$) and HybridEP (HT) with NCCL’s production infrastructure, NCCL EP delivers high-performance MoE communication within the standard GPU communication ecosystem.